OpenStudy (steve816):

I got a quick simple chemistry question and help would be awesome :) Calculate the number of atoms in a 2.76 x 10^3 g sample of calcium.

OpenStudy (baru):

u know mass number of calcium?

OpenStudy (steve816):

Yeah, the mass is 40.08 amu

OpenStudy (baru):

calcium: 20 protons and 20 neutrons mass number: 40

OpenStudy (baru):

do you know avagadro's number?

OpenStudy (steve816):

6.022 x 10^23

OpenStudy (baru):

40 gram of calcium contains 6.022 x 10^23 atoms can you tell how to find number of atoms in 1 gram?

OpenStudy (steve816):

Divide by 40?

OpenStudy (baru):

yea 6.022 x 10^23/40 is the number of atoms in 1 gram of calcium so the number of atoms in 2.76 x 10^3 grams is?

OpenStudy (steve816):

So the avagadro's number is the amount in grams?

OpenStudy (steve816):

And also, is the amu in grams too?

OpenStudy (baru):

u know how to proceed? no amu is a unit on its own. if mass number(protons +neutrons) of some element is M, then M grams of the substance contains "6.022 x 10^23" atoms in it.

OpenStudy (steve816):

Ohhh, okay.

OpenStudy (steve816):

So how would I enter it in my calculator?

OpenStudy (baru):

you will have to do the 10^ stuff by hand

OpenStudy (steve816):

Help please? Not good with scientific notation :(

OpenStudy (baru):

\[\frac{ 6.022 \times 10^{23} \times 2.76 \times 10^3 }{ 40 }\]

OpenStudy (baru):

\[\frac{ 6.022 \times 2.76 \times (10^3\times 10^{23}) }{ 40 }\]

OpenStudy (baru):

do everything outside the brackets by calculator

OpenStudy (steve816):

\[\frac{ 16.621 \times 10^{26}}{ 40 }\]

OpenStudy (anonymous):

Good